# Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$

### v1

@@ 6 Summary
We have stacked $\sim 100$k $K$-band selected COSMOS and UDS galaxies in the IR Herschel and JCMT bands in bins of stellar mass and redshift in order to investigate the SFR evolution out to $z\sim 7$. Inverse variance stacking was used, where, for the Herschel SPIRE bands, background sources were found to contribute significantly to the stacked fluxes. After removing the background sources contribution (fitted with the Gaussian function), the IR luminosities and hence star formation rates were determined and the functional forms of the SFR evolution with stellar mass and redshift found. This was done for both the star-forming and whole (star-forming and quiescent) galaxy samples, where the quiescent sources were selected using the $UVJ$ color-color space. The main findings of this work can be summarized as follows:
@@
    The SFR of star-forming galaxies populating the main sequence exhibits a nearly linear growth with stellar mass at any given redshift at low masses. Above the so-called bending mass, $M_{0}$, the main sequence flattens towards an asymptotic value of the SFR. At constant stellar mass, the SFR also increases with redshift – the redshift evolution of the normalization of the main sequence. This behavior is consistent with the scenario in which the evolution of the star-forming main sequence is driven by the position of the galaxy on the integrated Schmidt-Kennicutt relation, given its molecular mass fraction (e.g. Santini et al. 2014; Wang et al. 2022). We find that the redshift evolution of the bending mass, $M_{0}$, is exponential, increasing sharply from $z\sim 0.5$ to $z\sim 2$ and flattening at $z>2$. The similarity to the shape of the relation between the stellar mass, above which shocks can efficiently heat incoming baryons, and redshift, found in Daddi et al. (2022), suggests that the flattening of the main sequence is likely caused by the phasing out of the cold accretion.
@@
    We investigate the evolution of the peak dust temperature with IR luminosity and redshift, by fitting the stacked photometry with dust emission curve of Casey (2012). Similarly to Drew & Casey (2022), an increase of $T_{\rm d}$ with IR luminosity was found. However, when confined to any given redshift, our data does not show any significant evolution with $L_{\rm IR}$. We conclude that the increase of $T_{\rm d}$ with IR luminosity is driven by the evolution of the $L_{\rm IR}$ with redshift for the main-sequence galaxies. We find that the dust temperature increases linearly with redshift, consistent with the scenario in which higher values of $T_{\rm d}$ for star-forming galaxies are driven by the increase in their specific star formation rates (e.g. Liang et al. 2019).
@@
    The evolution of the main sequence found in this work was compared with the relations of Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023). We find that the normalization of our main sequence at $z\lesssim 2$ for masses below the bending mass, $M_{0}$, is slightly lower, while the SFRs above $M_{0}$ at high redshifts exhibit larger values (except for the MS of Speagle et al. 2014, where no main-sequence bending was found). We attribute these inconsistencies to the dust temperatures assumed, when fitting the IR data. As found in this work, $T_{\rm d}$ increases from $\sim 20$ K to $\sim 50$ K between redshifts 0 and 7 and hence, assuming dust temperatures at the level of $\sim 30$ K, as is often done in the absence of the far-IR data, will correspondingly affect the determined IR luminosities (and hence SFRs). We also note that the relatively higher values of the SFRs at high redshifts for $M_{\ast}>10^{11}\,{\rm M_{\odot}}$ can be, at least in part, caused by the contamination of the main-sequence sample with starburst galaxies.
@@
    Adopting different shapes of the SFR-$M_{\ast}$ relation, we simulate the time evolution for the observed galaxy stellar mass functions, GSMFs, of Davidzon et al. (2017) and McLeod et al. (2021). For the star-forming sample we find that the GSMF ‘grows’ too fast, producing simulated relations significantly above the observed ones. By performing the simulations using the SFR-$M_{\ast}$ relation, found using both star-forming and quiescent galaxies, we confirm that the overprediction of star-forming galaxies number densities is caused by a significant fraction of sources evolving into the quiescent phase and, thereby, leaving the star-forming main sequence. In addition, we find the simulated low-mass end of the GSMF to give higher number densities at high $z$, indicating that the low-mass slope of the main sequence likely evolves towards steeper values with redshift. However, due to the significant uncertainties, more detailed analysis based on larger high-$z$ samples is required.
@@ Acknowledgements.
Appendix A Plots & Tables
Figure 9: Example stacked images, shown in bins of stellar mass and redshift, for the Herschel SPIRE 500 ${\rm\mu m}$ band.
Figure 10: Stacked far-IR photometry for the whole sample of galaxies (star-forming and quiescent), listed in the top panel of Table 2. The best-fit dust emission curves, adopted from Casey (2012), are shown in solid green with $1\sigma$ standard deviation in grey.
Figure 11: As in Figure 10 but for star-forming sample only – bottom panel in Table 2.
@@ Appendix A Plots & Tables
$9.50\phantom{0}9.25\leq\mathcal{M}<\phantom{0}$ 9.50 $-0.13\pm 0.06$
$9.75\phantom{0}9.50\leq\mathcal{M}<\phantom{0}$ 9.75 $\phantom{-}0.18\pm 0.04$ $0.53\pm 0.05$
$10.00\phantom{0}9.75\leq\mathcal{M}<$ 10.00 $\phantom{-}0.37\pm 0.05$ $0.74\pm 0.05$ $1.04\pm 0.06$
$10.2510.00\leq\mathcal{M}<$ 10.25 $\phantom{-}0.54\pm 0.05$ $0.91\pm 0.05$ $1.21\pm 0.05$ $1.39\pm 0.06$ $1.52\pm 0.07$
$10.5010.25\leq\mathcal{M}<$ 10.50 $\phantom{-}0.60\pm 0.05$ $1.05\pm 0.05$ $1.38\pm 0.05$ $1.50\pm 0.06$ $1.68\pm 0.07$ $1.88\pm 0.07$
$10.7510.50\leq\mathcal{M}<$ 10.75 $\phantom{-}0.66\pm 0.05$ $1.14\pm 0.05$ $1.46\pm 0.05$ $1.74\pm 0.05$ $1.82\pm 0.06$ $1.98\pm 0.07$ $11.0010.50\leq\mathcal{M}<$ 11.00 $2.34\pm 0.08$ $2.47\pm 0.14$
$11.0010.75\leq\mathcal{M}<$ 11.00 $\phantom{-}0.69\pm 0.05$ $1.20\pm 0.05$ $1.59\pm 0.05$ $1.78\pm 0.05$ $1.96\pm 0.06$ $2.19\pm 0.07$
$11.2511.00\leq\mathcal{M}<$ 11.25 $\phantom{-}0.74\pm 0.05$ $1.22\pm 0.05$ $1.66\pm 0.05$ $1.98\pm 0.05$ $2.22\pm 0.06$ $2.42\pm 0.06$ $11.5011.00\leq\mathcal{M}<$ 11.50 $2.61\pm 0.07$ $2.80\pm 0.07$
$11.5011.25\leq\mathcal{M}<$ 11.50 $\phantom{-}0.60\pm 0.07$ $1.27\pm 0.05$ $1.71\pm 0.06$ $2.13\pm 0.06$ $2.40\pm 0.06$ $2.55\pm 0.07$
$9.50\phantom{0}9.25\leq\mathcal{M}<\phantom{0}$ 9.50 $-0.05\pm 0.06$
$9.75\phantom{0}9.50\leq\mathcal{M}<\phantom{0}$ 9.75 $\phantom{-}0.27\pm 0.04$ $0.57\pm 0.05$
$10.00\phantom{0}9.75\leq\mathcal{M}<$ 10.00 $\phantom{-}0.47\pm 0.05$ $0.80\pm 0.05$ $1.06\pm 0.06$
$10.2510.00\leq\mathcal{M}<$ 10.25 $\phantom{-}0.68\pm 0.05$ $1.01\pm 0.05$ $1.27\pm 0.06$ $1.42\pm 0.06$ $1.53\pm 0.08$
$10.5010.25\leq\mathcal{M}<$ 10.50 $\phantom{-}0.81\pm 0.05$ $1.22\pm 0.05$ $1.49\pm 0.05$ $1.59\pm 0.06$ $1.76\pm 0.06$ $1.89\pm 0.07$
$10.7510.50\leq\mathcal{M}<$ 10.75 $\phantom{-}0.93\pm 0.05$ $1.35\pm 0.05$ $1.63\pm 0.05$ $1.85\pm 0.05$ $1.93\pm 0.06$ $2.03\pm 0.07$ $11.0010.50\leq\mathcal{M}<$ 11.00 $2.38\pm 0.08$ $2.47\pm 0.14$
$11.0010.75\leq\mathcal{M}<$ 11.00 $\phantom{-}1.01\pm 0.05$ $1.47\pm 0.05$ $1.78\pm 0.05$ $1.93\pm 0.05$ $2.07\pm 0.06$ $2.27\pm 0.07$
$11.2511.00\leq\mathcal{M}<$ 11.25 $\phantom{-}1.12\pm 0.05$ $1.51\pm 0.05$ $1.88\pm 0.05$ $2.11\pm 0.05$ $2.31\pm 0.06$ $2.50\pm 0.06$ $11.5011.00\leq\mathcal{M}<$ 11.50 $2.64\pm 0.08$ $2.80\pm 0.07$
$11.5011.25\leq\mathcal{M}<$ 11.50 $\phantom{-}1.12\pm 0.06$ $1.65\pm 0.05$ $1.96\pm 0.05$ $2.27\pm 0.05$ $2.53\pm 0.06$ $2.61\pm 0.07$

### v2

@@ 6 Summary
We have stacked $\sim 100$k $K$-band selected COSMOS and UDS galaxies in the IR Herschel and JCMT bands in bins of stellar mass and redshift in order to investigate the SFR evolution out to $z\sim 7$. Inverse variance stacking was used, where, for the Herschel SPIRE bands, background sources were found to contribute significantly to the stacked fluxes. After removing the background sources’ contribution (fitted with the Gaussian function), the IR luminosities, and hence the SFRs, were determined and the functional forms of the SFR evolution with stellar mass and redshift found. This was done for both the star-forming and whole (star-forming and quiescent) galaxy samples, where the quiescent sources were selected using the $UVJ$ color-color space. The main findings of this work can be summarized as follows:
@@
    The SFR of SFGs populating the MS exhibits a nearly linear growth with stellar mass at any given redshift at low masses. Above the so-called bending mass, $M_{0}$, the MS flattens toward an asymptotic value of the SFR. At constant stellar mass, the SFR also increases with redshift – the redshift evolution of the normalization of the MS. This behavior is consistent with the scenario in which the evolution of the star-forming MS is driven by the position of the galaxy on the integrated SK relation, given its molecular mass fraction (e.g., Santini et al. 2014; Wang et al. 2022). We find that the redshift evolution of the bending mass, $M_{0}$, is exponential, increasing sharply from $z\sim 0.5$ to $z\sim 2$ and flattening at $z>2$. The similarity to the shape of the relation between the stellar mass, above which shocks can efficiently heat incoming baryons, and redshift, found in Daddi et al. (2022), suggests that the flattening of the MS is likely caused by the phasing out of the cold accretion.
@@
    We investigated the evolution of the peak dust temperature with IR luminosity and redshift, by fitting the stacked photometry with the dust emission curve of Casey (2012). Similarly to Drew & Casey (2022), an increase in $T_{\rm d}$ with IR luminosity was found. However, when confined to any given redshift, our data does not show any significant evolution with $L_{\rm IR}$. We conclude that the increase in $T_{\rm d}$ with IR luminosity is driven by the evolution of the $L_{\rm IR}$ with redshift for the MS galaxies. We find that the dust temperature increases quadratically with redshift, consistent with the scenario in which higher values of $T_{\rm d}$ for SFGs are driven by the increase in their sSFRs (e.g., Liang et al. 2019).
@@
    The evolution of the MS found in this work was compared with the relations of Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023). We find that the normalization of our MS at $z\lesssim 2$ for masses below the bending mass, $M_{0}$, is slightly lower, while the SFRs above $M_{0}$ at high redshifts exhibit larger values (except for the MS of Speagle et al. 2014, where no MS bending was found). We attribute these inconsistencies to the dust temperatures assumed, when fitting the IR data. As was found in this work, $T_{\rm d}$ increases from $\sim 20$ K to $\sim 60$ K between redshifts 0.5 and 5.5, and hence assuming dust temperatures at the level of $\sim 30$ K, as is often done in the absence of the FIR data, will correspondingly affect the determined IR luminosities (and hence the SFRs). We also note that the higher values of the SFRs at high redshifts for $M_{\ast}>10^{11}\,{\rm M_{\odot}}$ could be, at least in part, caused by the contamination of the MS sample by starburst galaxies.
@@
    Adopting different shapes of the SFR-$M_{\ast}$ relation, we simulated the time evolution for the observed GSMFs of Davidzon et al. (2017) and McLeod et al. (2021). For the star-forming sample, we find that the GSMF “grows” too fast, producing simulated relations significantly above the observed ones. By performing the simulations using the SFR-$M_{\ast}$ relation, found using both star-forming and quiescent galaxies, we confirm that the overprediction of SFG number densities is caused by a significant fraction of sources evolving into the quiescent phase and thereby leaving the star-forming MS. In addition, we find the simulated low-mass end of the GSMF gives higher number densities at high $z$, indicating that the low-mass slope of the MS likely evolves toward steeper values with redshift. However, due to the significant uncertainties, more detailed analysis based on larger high-$z$ samples is required.
7 Data availability
Example stacked image and best-fit SEDs are available here.
@@ Acknowledgements.
Appendix A Additional table
@@
$9.50\phantom{0}9.25\leq\mathcal{M}<\phantom{0}$ 9.50 $0.01\pm 0.06$ $0.20\pm 0.05$ $0.28\pm 0.06$ $0.51\pm 0.06$
$9.75\phantom{0}9.50\leq\mathcal{M}<\phantom{0}$ 9.75 $0.18\pm 0.05$ $0.42\pm 0.05$ $0.51\pm 0.05$ $0.68\pm 0.05$ $0.85\pm 0.05$
$10.00\phantom{0}9.75\leq\mathcal{M}<$ 10.00 $0.39\pm 0.05$ $0.54\pm 0.05$ $0.66\pm 0.05$ $0.84\pm 0.05$ $1.04\pm 0.06$ $1.21\pm 0.06$
$10.2510.00\leq\mathcal{M}<$ 10.25 $0.44\pm 0.05$ $0.65\pm 0.05$ $0.81\pm 0.05$ $0.97\pm 0.05$ $1.18\pm 0.05$ $1.33\pm 0.06$ $1.57\pm 0.06$
$10.5010.25\leq\mathcal{M}<$ 10.50 $0.55\pm 0.05$ $0.64\pm 0.06$ $0.86\pm 0.05$ $1.11\pm 0.05$ $1.28\pm 0.05$ $1.49\pm 0.05$ $1.68\pm 0.06$ $2.10\pm 0.08$
$10.7510.50\leq\mathcal{M}<$ 10.75 $0.47\pm 0.06$ $0.74\pm 0.05$ $0.90\pm 0.05$ $1.17\pm 0.05$ $1.41\pm 0.05$ $1.54\pm 0.05$ $1.80\pm 0.06$ $2.22\pm 0.09$
$11.0010.75\leq\mathcal{M}<$ 11.00 $0.70\pm 0.06$ $0.78\pm 0.06$ $0.85\pm 0.06$ $1.26\pm 0.05$ $1.44\pm 0.05$ $1.73\pm 0.05$ $2.01\pm 0.06$ $2.38\pm 0.09$
$11.2511.00\leq\mathcal{M}<$ 11.25 $0.61\pm 0.08$ $0.79\pm 0.07$ $0.92\pm 0.06$ $1.27\pm 0.06$ $1.55\pm 0.06$ $1.81\pm 0.06$ $2.20\pm 0.06$ $2.60\pm 0.08$
$11.5011.25\leq\mathcal{M}<$ 11.50 $1.02\pm 0.08$ $1.53\pm 0.08$ $2.06\pm 0.07$ $2.31\pm 0.09$ $2.71\pm 0.10$
$9.50\phantom{0}9.25\leq\mathcal{M}<\phantom{0}$ 9.50 $0.05\pm 0.05$ $0.26\pm 0.06$ $0.35\pm 0.07$ $0.52\pm 0.06$
$9.75\phantom{0}9.50\leq\mathcal{M}<\phantom{0}$ 9.75 $0.25\pm 0.05$ $0.49\pm 0.05$ $0.57\pm 0.05$ $0.74\pm 0.05$ $0.87\pm 0.06$
$10.00\phantom{0}9.75\leq\mathcal{M}<$ 10.00 $0.50\pm 0.05$ $0.69\pm 0.06$ $0.81\pm 0.06$ $0.91\pm 0.06$ $1.12\pm 0.06$ $1.24\pm 0.06$
$10.2510.00\leq\mathcal{M}<$ 10.25 $0.62\pm 0.06$ $0.84\pm 0.06$ $1.00\pm 0.05$ $1.10\pm 0.05$ $1.28\pm 0.06$ $1.43\pm 0.06$ $1.61\pm 0.06$
$10.5010.25\leq\mathcal{M}<$ 10.50 $0.76\pm 0.05$ $0.92\pm 0.06$ $1.12\pm 0.05$ $1.28\pm 0.05$ $1.44\pm 0.05$ $1.62\pm 0.05$ $1.73\pm 0.06$ $2.16\pm 0.08$
$10.7510.50\leq\mathcal{M}<$ 10.75 $0.77\pm 0.06$ $1.06\pm 0.06$ $1.24\pm 0.05$ $1.39\pm 0.05$ $1.57\pm 0.05$ $1.70\pm 0.05$ $1.90\pm 0.06$ $2.26\pm 0.10$
$11.0010.75\leq\mathcal{M}<$ 11.00 $1.01\pm 0.06$ $1.11\pm 0.06$ $1.25\pm 0.06$ $1.49\pm 0.06$ $1.63\pm 0.06$ $1.88\pm 0.06$ $2.08\pm 0.06$ $2.38\pm 0.09$
$11.2511.00\leq\mathcal{M}<$ 11.25 $0.96\pm 0.11$ $1.17\pm 0.09$ $1.42\pm 0.08$ $1.52\pm 0.07$ $1.79\pm 0.07$ $1.96\pm 0.06$ $2.28\pm 0.06$ $2.62\pm 0.08$
$11.5011.25\leq\mathcal{M}<$ 11.50 $1.14\pm 0.32$ $1.39\pm 0.14$ $1.73\pm 0.10$ $2.27\pm 0.08$ $2.42\pm 0.10$ $2.74\pm 0.09$